OpenStudy (anonymous):

A boy can row a boat at a constant rate of 5 mi/hr in still water. He rows upstream for 15 minutes and downstream, returning to his starting point in another 12 minutes. What is the rate of the current? What is the total distance traveled?

OpenStudy (anonymous):

15/60 = D/(5-C) 12/60 = D/(5+C) C = CURRENT RATE D = 1/2 TOTAL DISTANCE

OpenStudy (anonymous):

Rate * Time = Distance. Solve the following for c, the speed of the current in miles/hour:\[(5-c)\frac{15}{60}\text{ = }(5+c)\frac{12}{60}\]\[\left\{c\to \frac{5}{9}\right\} \]c = 5/9 mph\[\text{distance} =\left(5-\frac{5}{9}\right)\frac{15}{60}=\frac{10}{9}=1.11\text{miles} \]